S2 (use the end Solve shifting theoren 11, ostat the problem y-y -2y = f(t), f(t)= {0, t71 y (ol = 1, y(o)=0
Answer #1

Eniven the differential equahon. Solution 12, Ost <1 y - 4 - 2y = f(t), where To, txt -(1) Taking the flt) y(o)= 1, Y0=0 the--<!{}}+{{}+ {} Exit it theat 1 + { se)} = - + 3 et est and ers th -} thwrem. 8(828-2) Note: if x{f(s)} = f(t) then tlfe-a